OpenStudy (anonymous):

Which of the following sequences are geometric? A. -4, -2, -1, -0.5, -0.25, -0.125 B. 2, 3, 5, 8, 13, 21 C. 2, 5, 8, 11, 14, 17 D. 6, 18, 54, 162, 486

OpenStudy (ranga):

See if they have a common ratio. Divide first term by second term. Is that the same when second term is divide by third term? Third by fourth? If yes, then it is a geometric sequence. If not, then it is not.

OpenStudy (anonymous):

okay, so D is a possible answer right?

OpenStudy (ranga):

You can reverse the order of division too. 2nd term / 1st term = 3rd term / 2nd term = 4th term / 3rd term = r Or 1st term / 2nd term = 2n term / 3rd term = 3rd term / 4th term = r

OpenStudy (ranga):

Yes, D is a geometric sequence.

OpenStudy (anonymous):

doing the math it doesnt seem like the other 3 are possible answers. am i missing something or is that correct?

OpenStudy (ranga):

Try A again.

OpenStudy (anonymous):

Oh, would A be correct because it increases by .25 each time?

OpenStudy (ranga):

not "increases by .25" The common ratio is 1/2. Just multiply each term by 1/2 and you will get the next term.

OpenStudy (ranga):

Geometric Sequence goes like: a, ar, ar^2, ar^3, ar^4, ......

OpenStudy (kc_kennylau):

Where \(a\) is the starting term and \(r\) is the common ratio

OpenStudy (ranga):

In case A) a= -4 and r = 1/2

OpenStudy (kc_kennylau):

and in case D) a=6 and r=3

OpenStudy (ranga):

You multiply each term by the same constant (called the common ratio) and you will get the next term.

OpenStudy (ranga):

case D) 6 6*3 = 18 18*3 = 54 54*3 = 162 162*3 = 486

OpenStudy (anonymous):

oh that completely makes sense! for some reason I was thinking of something totally different. thank you both! :)

OpenStudy (ranga):

you are welcome.

OpenStudy (kc_kennylau):

no problem :)